# CRoP: Context-wise Robust Static Human-Sensing Personalization

### v4

@@ 2.3. Test-Time Adaptation (TTA)
Continual learning requires repeated training cycles to adapt models to new information, demanding high computational resources, storage, and long-term data retention, which raises privacy concerns (Verma et al., 2023). This makes traditional continual learning impractical for sensitive domains like healthcare (Harun et al., 2023).
Test-time adaptation offers a more efficient and felxible alternative, allowing models to adapt in real-time during deployment by making adjustments based on incoming data without necessitating a complete retraining process. By focusing on adapting the model in real-time, test-time adaptation minimizes resource consumption as compared to continual learning approaches. Wang et al. (2022a) relies on generating pseudo-labels using weighted averaging and augmentation-averaged predictions, which can introduce bias (Liang et al., 2020). While Wu et al. (2024) achieves test-time domain adaptation by manipulating batch-normalization layers, thus enforcing a restriction on model architecture. However, Continual Test Time Domain Adaptation (CoTTA) (Wang et al., 2022b) does not have such restrictions and allows the use of off-the-shelf models.
CoTTA (Wang et al., 2022b) is an unsupervised learning approach designed to enhance model adaptation in dynamic environments. It utilizes a teacher model, which is initially set up as a replica of the generic model, to generate pseudo-labels for training. During each iteration, the teacher model is updated using current model state through a weighted sum, allowing the model to effectively account for evolving data patterns over time. Additionally, the current model state is subject to a stochastic restoration of weights, implemented through a Bernoulli distribution with a very low probability of success. This randomization introduces an element of variability that helps prevent overfitting and encourages exploration of the weight space, further enhancing the model’s ability to generalize to unseen data.
Since CoTTA addresses a problem setting similar to ours, we adopt it as one of our baseline approaches. In our setup, data from the available context can be leveraged for test-time adaptation, while unseen context data is reserved solely for empirical evaluation. However, test-time adaptation methods, including CoTTA, tend to increase resource demands during inference, leading to potential delays that can impact real-time usability. This additional computational overhead highlights a key limitation of test-time adaptation in applications where real-time response is essential.
@@ 2.4. Personalization Approaches
A few static personalization approaches (Burns et al., 2022; Duan et al., 2023; Liu et al., 2022) aimed for the additional goal of out-of-distribution robustness. However, these methods require access to the generic model—either to make specific design choices (Burns et al., 2022), which prevents them from utilizing off-the-shelf models, or to incorporate knowledge about the target user’s data distribution during the generic model’s training phase (Duan et al., 2023; Liu et al., 2022), raising privacy concerns, particularly in sensitive clinical applications. These requirements do not align with the research objectives of this paper, making them unsuitable as baselines.
For instance, EMGSense (Duan et al., 2023) is concerned with robustness to distributional shifts in Surface Electromyography (EMG) signals caused by the heterogeneity of biological factors across users. The generic model consists of a multi-model voting ensemble DNN that needs to be trained on N source distributions and some unlabeled data from target users. To learn user-specific features, this model undergoes multiple training iterations, each using updated user-specific data.
Also, MobilePhys (Liu et al., 2022) addresses distributional shifts that can arise due to environmental conditions (e.g., lighting, motion) as well as individual physiological and visual (e.g., clothing, posture) differences in camera-based contactless Photoplethysmography. The generic model is trained using meta-learning, which uses data from target users to enable quick adaptation to distributional shifts during personalization.
Consequently, the major drawback faced by EMGSense(Duan et al., 2023) and MobilePhys (Liu et al., 2022) is the requirement of knowledge of the target user’s data in order to train the generic model. This necessitates the transfer of the user’s unlabelled data from the user’s device to a central server. Such data can contain sensitive information, and thus, its transfer suffers from legal and regulatory barriers (Food and Administration, 2019). More importantly, since the training of the generic model requires a subset of each target user’s unlabelled data, the introduction of any new user will require re-training of the generic model and its distribution, which is an overhead.
Recently, Burns et al. (2022) extends the idea of triplet loss (Schroff et al., 2015) to the task of personalization (PTN). The optimization objective combines the minimization of the Euclidean distance between data from the same target classes while maximizing the Euclidean distance between different target classes in order to learn an embedding optimized for the desired task. The features extracted using DNN have been shown to be superior to the engineered features (Sani et al., 2017). Thus, Burns et al. (2022) uses these features to train a KNN for the final prediction task. The approach has been shown to be robust for out-of-distribution data.
@@ 2.5. Selecting Baselines
Feature $\rightarrow$ Off-the-shelf Generic Data Adaptation to Robustness to
Approach $\downarrow$ Generic model unseen available context unseen Context
Arjovsky et al. (2020); Ganin et al. (2016) -NA-
Long et al. (2016)
Li et al. (2021); Ganin and Lempitsky (2015)
SHOT (Liang et al., 2020)
Packnet(Mallya and Lazebnik, 2018) and Piggyback(Mallya et al., 2018)
CoTTA (Wang et al., 2022b)
EMGSense (Duan et al., 2023) and MobilePhys (Liu et al., 2022)
PTN (Burns et al., 2022)
CRoP
Table 1. Summary of Literature review
The problem addressed in this work has the following requirements: using an Off-the-shelf generic model, having no access to the generic data, and requiring adaptation to available context data and robustness to unseen data. As discussed in the previous sections and summarized in Table 1, the approaches SHOT (Liang et al., 2020), PTN (Burns et al., 2022), PackNet (Mallya and Lazebnik, 2018), Piggyback (Mallya et al., 2018), and Continual Test Time Domain Adaptation (CoTTA) (Wang et al., 2022b) comply with the first three requirements. Thus, we adapt these approaches to the problem statement addressed in this work and use them as our baselines. These approaches typically rely on training data from a specific context to adapt to that context. However, the scenario addressed in this work limits training data to only one context, with the expectation that the approach will also perform well in unseen contexts. Since data from unseen contexts is unavailable, we prevent these approaches from using it during training. Instead, the data used for training or adaptation is restricted to data belonging to the available context only, while the unseen context data is used merely for testing.
@@ 3. Motivation
Table 2. Performance Comparison in terms of inference accuracy of Generic model with conventionally trained personalized model
Table 2(a) compares the performance of generic and conventionally-finetuned(Hong et al., 2016) personalized models on each user’s data belonging to context C1 and evaluated to the same user’s disjoint data in both C1 (available) and C2 (unseen) contexts. It can be observed that conventional finetuning introduces a significant gain of $26.82\%$ for context C1’s data but at the cost of $24.63\%$ reduction in context C2. Similar patterns are seen when personalization is performed on context C2 as shown in Table 2(b). Thus, conventional finetuning-based personalization of the models using the limited data from one context can significantly worsen the model’s performance in an unseen context.
(a) Context C1
(b) Context C2
Figure 3. Heat map for the absolute magnitude of parameters belonging to penultimate layer for LeNet models finetuned using data from context (a) C1 and (b) C2
To investigate this discrepancy in performance, we compare the distribution of parameter magnitudes of the models personalized on contexts C1 and C2 using conventional finetuning, as shown using heat maps in Figures 3 (a) and 3 (b). Notably, there is a substantial difference in parameter magnitudes between models trained in different contexts. Additionally, the parameters represented by black pixels in Figure 3 have magnitudes close to zero, indicating two crucial aspects: (a) Their contribution to model inference is negligible, implying redundancy. (b) Interestingly, some of these parameters have high magnitudes in the personalized model of another context, indicating that parameters considered unimportant in one context may be crucial in another.
@@ 5.1. Rationale for The CRoP Approach Design
Model pruning is based on the idea that neural networks include redundant parameters, and removing these parameters minimally impacts the model’s performance (Luo et al., 2017; Zhu and Gupta, 2017). Consequently, pruning the fine-tuned personalized model ensures the retention of essential parameters to maintain accuracy for context $\mathcal{C}_{a}$, meaning, capturing the target user-specific traits. However, the pruned parameters can be replaced with corresponding parameters from the generic model, effectively restoring generic knowledge learned across all contexts on those generic model parameters. This restoration may enhance generalizability, ensuring robust performance in unseen contexts $\mathcal{C}_{u}$. The approach presented in this paper is founded on this insightful strategy.
@@ 5.2. CRoP Approach
$\mathcal{M}^{P^{a}_{i}}_{\theta^{\downarrow}}=\text{ToleratedPrune}(\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}},\tau)$
4:  Copy the parameters of generic models to the pruned parameters in the personalized pruned model,
@@
The approach initiates by finetuning the generic model $\mathcal{M}^{G}_{\theta}$ on data $\mathcal{D}^{a}_{i}$, concurrently applying $\ell_{1}$ regularization to penalize model parameters (line 2). This regularization encourages sparsity by specifically targeting the magnitude of redundant parameters (Mayank, 2023). This step is followed by the pruning of redundant weights using the ‘ToleratedPrune’ module (line 3). The pruned weights are then replaced by the corresponding weights from the generic model $\mathcal{M}_{G}$ (line 4) to restore generalizability; this hybrid model is referred to as the ‘Mixed Model.’ This step leads to the modification of the activated paths in the personalized model, resulting in changes in the model inferences. However, since the newly activated paths are determined by weights retained from two models and not learned from data patterns, there is a consequent loss of accuracy, as shown and discussed in Section 7.2. To mitigate such a loss, as a final step, the Mixed Model undergoes fine-tuning once again on the data from the available context $\mathcal{D}^{a}_{i}$ (line 5). The detailed explanation of each of these steps is as follows:
@@
    Minimal loss in $\mathcal{C}_{a}$ accuracy: A high fraction of parameters have close to zero magnitudes, and their removal results in minimal information loss for context $\mathcal{C}_{a}$; thus, the adverse impact of pruning in context $\mathcal{C}_{a}$ is minimized.
@@
    Maximal generalization: The inclusion of regularization aids ToleratedPrune (discussed below) module in efficiently pruning a higher number of parameters, which are then replaced with weights from the generic model. This restores information from the generic model, contributing to enhanced accuracy in unseen contexts.
@@
ToleratedPrune Module (Algorithm 1 – Step $3$):
Algorithm 2 outlines the ToleratedPrune module, taking a model $\mathcal{M}_{\theta}$, pruning tolerance $\tau$, and the dataset $\mathcal{D}$ as inputs. It initiates with a modest pruning amount of $k$ and incrementally increases this amount by $k^{\prime}$ until the model’s accuracy exhibits a drop of $\tau$ percent on $\mathcal{D}$. Here, $k$ and $k^{\prime}$ are hyperparameters within the range of $(0,1)$. The module returns $\mathcal{M}_{\theta^{\downarrow}}$, representing the pruned state of the model before the last pruning iteration. This state is such that further pruning would result in a higher accuracy loss on dataset $\mathcal{D}$ than the tolerable amount $\tau$. This module performs pruning leveraging the conventional magnitude-based unstructured pruning (Zhu and Gupta, 2017).
@@
Notably, model pruning is often followed by a finetuning step (Luo et al., 2017; Zhu and Gupta, 2017; Liu et al., 2020), where the pruned model undergoes re-training to recover the performance lost during the pruning process. We have observed that, despite the Mixed Model exhibiting improved performance in the unseen context, there is a notable loss of accuracy in the available context due to inconsistent activated paths, as discussed earlier. Thus, the resulting Mixed Model is fine-tuned using the available data $\mathcal{D}^{a}_{i}$. Goyal et al. (2023) suggests that fine-tuning process should mirror pre-training for effective generalization. Therefore, our fine-tuning objective aligns with the pre-training objective used in Line 2 for optimal results.
During finetuning, the model state, including the mixed model, with the best validation loss on the seen context, is selected. We found that for some individuals, the mixed model is chosen as the optimal model, which indicates that for some individuals, further finetuning is not required, and our approach can automatically handle that scenario.
@@ 6. Experiments
This section displays the empirical efficacy of the presented approach. Section 6.1 provides a detailed discussion of the four human sensing datasets and their pre-processing used in our evaluations: PERCEPT-R (Benway et al., 2023), WIDAR (Zhang et al., 2022), ExtraSensory (Vaizman et al., 2017) and a Stress-sensing dataset (Xiao et al., 2024), and also explains the corresponding model architectures as used in literature. Section 6.2 details the metrics of evaluation used in this work along with their practical relevance. This is followed by a detailed discussion of the empirical comparison of CRoP with five baselines SHOT(Liang et al., 2020), Packnet(Mallya and Lazebnik, 2018), Piggyback(Mallya et al., 2018), CoTTa(Wang et al., 2022b) and PTN(Burns et al., 2022) in Section 6.3. Additionally, detailed discussions about interesting patterns, such as the impact of generic model quality on personalization, etc., are also provided.
@@ 6.1. Datasets and models
The sound / r / has been recognized as the most frequently impacted sound in residual speech sound disorders in American English (Lewis et al., 2015) and considered to be the most difficult sound to treat. The PERCEPT-R Corpus was collected during 34 different cross-sectional and longitudinal studies of speech (Benway et al., 2023) for automated speech analysis of / r /. The data used in this study come from the prospectively collected (Benway and Preston, 2023), and corpus version 2.2.2, which includes both the publicly available open access subset (2.2.2p) and privately held data that was not published in the open access subset after a review of consent/assent permissions. Items in the PERCEPT-R Corpus v2.2.2 primarily consist of single-word speech audio collected during clinical trials involving children with speech sound disorders affecting / r /, along with age-matched peers with typical speech. The full corpus contains 179,076 labeled utterances representing 662 single-rhotic words and phrases. Each audio file is paired with a ground-truth label representing listener judgments of rhoticity, derived by averaging binary ratings (0 = derhotic, 1 = fully rhotic) from multiple listeners. For this study, the heuristic threshold for converting these averaged ratings into binary ground-truth labels was 0.66.
@@
To use this dataset for our personalization evaluation, we collaborated with clinical experts to identify and acquire annotations of $16$ participants who had correct and incorrect pronunciations of / r / sound at pre-treatment (baseline-phase) and during different treatment phases. As outlined in Section 4, the evaluation treats pre-treatment data as available context data, while data from other treatment phases serve as unseen context data.
WIDAR:
@@
WIDAR is a dataset collected for the purpose of gesture recognition. It was collected using off-the-shelf WiFi links (one transmitter and at least 3 receivers). 17 users performed 15 different gestures at different rooms and orientations (of the person). The channel state information is collected from these devices with amplitude noises and phase offsets removed as a preprocessing step. The two contexts used for the current work are decided based on the orientations of the torso data and room ID. Room 1 is a classroom with a number of objects (e.g., desks and chairs) in it, and Room 2 is a nearly empty hallway. The dissimilar data distributions can be attributed to the differences in the amount of interruptions in WiFi signals. We followed the same normalization methods as Yang et al. (2023).
ExtraSensory:
@@
ExtraSensory is a human activity recognition dataset collected using the ExtraSensory mobile application. A number of features were collected from different cellular devices and smartwatches, though we just used the accelerometer and gyroscope features obtained from the cellular devices. Labels for activities were self-reported by the users through the mobile application. For our evaluations on ExtraSensory, 5 users were left out of training a single generic model. The contexts are decided based on the location of the phone: hand, pocket, and bag.
Stress Sensing Dataset:
@@
This dataset measures the physiological impacts of various kinds of Stress. The dataset is collected using Empatica E4 Wristband to extract features such as EDA (Electrodermal Activity), a skin temperature sensor ($4$ Hz), etc, contributing to a total of 34 features. The data was collected from 30 participants with different demographics, who were assigned the labels ‘Stressed’ or ‘Calm’ based on the different stress-inducing or calming tasks they experienced. In order to adapt this dataset for the problem addressed in this work, we collaborated with the original authors to identify participants who wore wristbands on both hands, which were then chosen for the task of personalization. Additionally, during the data collection, participants were asked to perform several activities, a few of which restricted participants’ movement while others allowed them to move. Based on this, the data was annotated with movement patterns, that is, still vs. moving. This additional annotation of the dataset will be made public along with this work. During this dataset’s evaluation, the context is defined by a combination of the hand on which the wristband was worn and whether or not the person was moving during the data collection.

### v5

@@ 2.3. Test-Time Adaptation (TTA)
Test-time adaptation offers a more efficient and flexible alternative, allowing models to adapt in real-time during deployment by making adjustments based on incoming data without necessitating a complete retraining process. By focusing on adapting the model in real-time, test-time adaptation minimizes resource consumption as compared to continual learning approaches. While Wu et al. (2024) achieves test-time domain adaptation by manipulating batch-normalization layers, thus enforcing a restriction on model architecture. However, Continual Test Time Domain Adaptation (CoTTA) (Wang et al., 2022) does not have such restrictions and allows the use of off-the-shelf models.
CoTTA (Wang et al., 2022) is an unsupervised learning approach designed to enhance model adaptation in dynamic environments. It utilizes a teacher model, which is initially set up as a replica of the generic model, to generate pseudo-labels for training. During each iteration, the teacher model is updated using weighted summation of current states of the teacher and the student model. This allows the teacher model to generate pseudo-lables which align with the evolving data patterns over time. Additionally, the current model state is subject to a stochastic restoration of weights, implemented through a Bernoulli distribution with a very low probability of success. This randomization introduces an element of variability that helps prevent overfitting and encourages exploration of the weight space, further enhancing the model’s ability to generalize to unseen data.
Figure 2. Model Structure for the Lenet Model used for Widar Dataset. The parameters of the highlighted layer are used for preliminary analysis
We adopt CoTTA as a baseline due to its similarity to our problem setting, where available context data supports test-time adaptation, and unseen context data is reserved for evaluation. However, test-time adaptation methods like CoTTA increase inference resource demands, potentially causing delays that hinder real-time usability—a key limitation in time-sensitive applications.
@@ 2.4. Personalization Approaches
A few static personalization approaches (Burns et al., 2022; Duan et al., 2023; Liu et al., 2022) aimed for the additional goal of out-of-distribution robustness. However, these methods require access to the generic model—either to make specific design choices (Burns et al., 2022), which prevents them from utilizing off-the-shelf models, or to incorporate knowledge about the target user’s data distribution during the generic model’s training phase (Duan et al., 2023; Liu et al., 2022), raising privacy concerns.
Recently, Burns et al. (2022) extends the idea of triplet loss (Schroff et al., 2015) to personalization (PTN). The optimization objective combines the minimization of the Euclidean distance between data from the same target classes while maximizing the Euclidean distance between different target classes in order to learn an embedding optimized for the desired task. The features extracted using DNN have been shown to be superior to the engineered features (Sani et al., 2017). Thus, Burns et al. (2022) uses these features to train a KNN for the final prediction task. The approach has been shown to be robust for out-of-distribution data.
@@ 2.5. Selecting Baselines
The problem addressed in this work has the following requirements: using an Off-the-shelf generic model, having no access to the generic data, and requiring adaptation to available context data and robustness to unseen data. As discussed in the previous sections, the approaches SHOT (Liang et al., 2020), PTN (Burns et al., 2022), PackNet (Mallya and Lazebnik, 2018), Piggyback (Mallya et al., 2018), and Continual Test Time Domain Adaptation (CoTTA) (Wang et al., 2022) comply with the first three requirements. Thus, we adapt these approaches to the problem statement addressed in this work and use them as our baselines. These approaches typically rely on training data from a specific context to adapt to that context. However, the scenario addressed in this work limits training data to only one context, with the expectation that the approach will also perform well in unseen contexts. Since data from unseen contexts is unavailable, we prevent these approaches from using it during training. Instead, the data used for training or adaptation is restricted to data belonging to the available context only, while the unseen context data is used merely for testing.
(a) Context C1
(b) Context C2
Figure 3. Heat map for the absolute magnitude of parameters belonging to penultimate layer for LeNet models finetuned using data from context (a) C1 and (b) C2. The first linear layer after the convolution layer has been used for this analysis. The size of the original weight matrix of the layer is 1536 X 128. However, we reshape heat map to 384 X 512 for better observability.
@@ 3. Motivation
Table 1. Inference accuracy comparison of generic model with conventionally trained personalized model
Table 4(a) compares the performance of generic and conventionally-finetuned (Hong et al., 2016) personalized models on each user’s data belonging to context C1 and evaluated to the same user’s disjoint data in both C1 (available) and C2 (unseen) contexts. It can be observed that conventional finetuning introduces a significant gain of $26.82\%$ for context C1’s data but at the cost of $24.63\%$ reduction in context C2. Similar patterns are seen when personalization is performed on context C2 as shown in Table 4(b). Thus, conventional finetuning-based personalization using limited data from one context can significantly worsen the model’s performance in an unseen context.
To investigate this discrepancy in performance, we compare the distribution of parameter magnitudes of the models personalized on contexts C1 and C2 using conventional finetuning, as shown using heat maps in Figures 3 (a) and 3 (b). Notably, both models started from the same generic model but were personalized independently, leading to different parameter magnitudes tailored to the traits of two different contexts. The structure of the Lenet Model used for the Widar Dataset is provided in Figure 2. The first linear layer after the convolution layers, that is, ‘fc1’, is chosen for this analysis as it captures the relative importance of input features (Goodfellow et al., 2016). Although the original size of the layer is 1536 X 128, we resized the corresponding heatmap to 384 X 512 to improve the visualization. In the heatmaps, the brighter shades of red represent a higher magnitude, while the black pixel represents a weight magnitude almost equal to zero. Thus, if a feature is considered less important, all corresponding parameters (connections) will have lower magnitudes, appearing as black rows in the heat map. Notably, there is a substantial difference in parameter magnitudes (i.e., color differences) between models trained in different contexts. Specifically, many rows in Figure 3(a) have black pixels, representing less useful features. These pixels represent parameters with magnitudes close to zero (empirically identified as lower than $1X10^{-4}$), and they comprise $18.66\%$ of the first linear layer for the model personalized on context C1. However, the average magnitude of the same parameters in the model finetuned for Context C2 is $0.0091$ (shown in Figure 3(b) with brighter red color pixels), which is approximately 100 times higher. This indicates two crucial aspects: (a) these parameters contribute minimally to model inference in Context C1, indicating redundancy for C1. (b) same parameters’ higher magnitudes in the model personalized in another context, C2, indicates that parameters considered unimportant in one context may be crucial in another.
@@ 5.1. Rationale for The CRoP Approach Design
Model pruning is based on the idea that neural networks include redundant parameters, and removing these parameters minimally impacts the model’s performance (Luo et al., 2017; Zhu and Gupta, 2017). Consequently, pruning the fine-tuned personalized model ensures the retention of essential parameters to maintain accuracy for context $\mathcal{C}_{a}$, meaning capturing the target user-specific traits present in context $\mathcal{C}_{a}$ in a substructure of the personalized model. Consequently, the values of the pruned parameters (that are zeroed out) can be replaced with corresponding parameter values from the generic model, effectively restoring generic knowledge learned across all contexts through those parameters of the generic model. This restoration may enhance generalizability, ensuring robust performance in unseen contexts $\mathcal{C}_{u}$. The approach presented in this paper is founded on this insightful strategy.
@@ 5.2. CRoP Approach
$\mathcal{M}^{P^{a}_{i}}_{\theta^{\downarrow}}=\text{ToleratedPrune}(\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}},\tau),$
where $\theta^{\downarrow}$ signifies that a subset of model parameters are non-zero.
4:  Copy the parameters of generic model to the pruned parameters in the personalized pruned model,
@@
The approach initiates by finetuning the generic model $\mathcal{M}^{G}_{\theta}$ on data $\mathcal{D}^{a}_{i}$, concurrently applying $\ell_{1}$ regularization to penalize model parameters (line 2). This regularization encourages sparsity by specifically targeting the magnitude of redundant parameters (Mayank, 2023). This step is followed by the pruning of redundant weights using the ‘ToleratedPrune’ module (line 3). The pruned weights are then replaced by the corresponding weights from the generic model $\mathcal{M}_{G}$ (line 4) to restore generalizability; this hybrid model is referred to as the ‘Mixed Model.’ This step can lead to model inconsistencies resulting in performance loss (Yadav et al., 2023; Daheim et al., 2024) for some users. To mitigate such a loss for the affected users, as a final step, the Mixed Model undergoes fine-tuning once again on the data from the available context $\mathcal{D}^{a}_{i}$ (line 5). The detailed explanation of each of these steps is as follows:
@@
    Minimal loss in $\mathcal{C}_{a}$ accuracy: A high fraction of parameters have close to zero, and their removal results in minimal information loss for context $\mathcal{C}_{a}$; thus, the adverse impact of pruning in context $\mathcal{C}_{a}$ is minimized.
@@
    Maximal generalization: The inclusion of regularization aids ToleratedPrune (discussed below) module in efficiently pruning a higher number of parameters, which are then replaced with weights from the generic model. This restores information from the generic model, contributing to enhanced accuracy in context $\mathcal{C}_{u}$.
ToleratedPrune Module (Algorithm 1 – Step $3$):
Algorithm 2 outlines the ToleratedPrune module, taking a model $\mathcal{M}_{\theta}$, pruning tolerance $\tau$, and the dataset $\mathcal{D}$ as inputs. Different individuals may exhibit varying degrees of deviation in their user-specific traits (Gong et al., 2023) compared to the generic traits learned by the generic model. As a result, fine-tuning a generic model requires different amounts of training to capture user-specific traits for each individual. Additionally, the model substructure necessary for learning these traits is shaped by regularization (Algorithm 1 Line 2) applied during the initial fine-tuning step. The combined effects of data heterogeneity, varying degrees of learning, and regularization lead to the necessity of requiring distinct substructures for each individual. Thus, choosing a single pruning amount for all individuals might not be the optimal pruning strategy.
@@
To solve this, the ToleratedPrune module iteratively computes an optimal sub-network for each individual’s available context $\mathcal{C}_{a}$. It initiates with a modest pruning amount of $k$ and incrementally increases this amount by $k^{\prime}$ until the model’s accuracy exhibits a drop of $\tau$ percent on $\mathcal{D}$. Here, $k$ and $k^{\prime}$ are hyperparameters within the range of $(0,1)$. The module returns $\mathcal{M}_{\theta^{\downarrow}}$, representing the pruned state of the model before the last pruning iteration. The symbol $\theta^{\downarrow}$ suggests that only a subset of this model’s parameters are non-zero. This state is such that further pruning would result in a higher accuracy loss on dataset $\mathcal{D}$ than the tolerable amount $\tau$. This module performs pruning leveraging the conventional magnitude-based unstructured pruning (Zhu and Gupta, 2017).
@@
Recent works have highlighted performance degradation while merging two trained models, attributing the loss to either parameter sign flip (Yadav et al., 2023) or change in gradient direction (Daheim et al., 2024). Both of these works show that directly merging two trained models can result in inconsistencies in the model, which may cause accuracy degradation, necessitating the need for a mechanism to address this. This is in line with our observation that despite the Mixed Model exhibiting improved performance in the unseen context, there is a notable loss of accuracy in the available context for some users. Lu et al. (2024) have shown that finetuning after merging different models can result in a performance boost. Thus, we include a final finetuning step in our personalization mechanism for the affected users. Goyal et al. (2023) suggests that fine-tuning process should mirror pre-training for effective generalization. Therefore, our final fine-tuning objective aligns with the pre-training objective used in Line 2 for optimal results.
During finetuning, the model state, including the mixed model, with the best validation loss on the available context, is selected. We found that for some individuals, the mixed model is chosen as the optimal model, which indicates that for some individuals, further finetuning is not required, and our approach can automatically handle that scenario.
@@ 6. Experiments
This section displays the empirical efficacy of the presented approach. Section 6.1 provides a detailed discussion of the four human sensing datasets and their pre-processing used in our evaluations: PERCEPT-R (Benway et al., 2023), WIDAR (Zhang et al., 2022), ExtraSensory (Vaizman et al., 2017) and a Stress-sensing dataset (Xiao et al., 2024), and also explains the corresponding model architectures as used in literature. Section 6.2 details the metrics of evaluation used in this work along with their practical relevance. This is followed by a detailed discussion of the empirical comparison of CRoP with five baselines SHOT(Liang et al., 2020), Packnet(Mallya and Lazebnik, 2018), Piggyback(Mallya et al., 2018), CoTTa(Wang et al., 2022) and PTN(Burns et al., 2022) in Section 6.3. Additionally, detailed discussions about interesting patterns, such as the impact of generic model quality on personalization, etc., are also provided.
@@ 6.1. Datasets and models
PERCEPT-R: The data used in this study come from corpus version v2.2.2 (Benway and Preston, 2023), which consist of single-word speech audio collected during clinical trials involving children with speech sound disorders affecting / r /, along with age-matched peers with typical speech. The full corpus contains 179,076 labeled utterances representing 662 single-rhotic words and phrases. Each audio file is paired with a ground-truth label representing listener judgments of rhoticity, derived by averaging binary ratings (0 = derhotic, 1 = fully rhotic) from multiple listeners.
@@
To use this dataset for our personalization evaluation, we collaborated with clinical experts to identify and acquire annotations of $16$ participants who had correct and incorrect pronunciations of / r / sound at pre-treatment (baseline-phase) and during different treatment phases. The evaluation treats pre-treatment data as available context data, while data from other treatment phases serve as unseen context data.
@@
WIDAR: WIDAR is a gesture recognition dataset collected using off-the-shelf WiFi links (one transmitter and at least three receivers). 17 users performed 15 gestures across different rooms and torso orientations. The channel state information is collected from these devices with amplitude noises and phase offsets removed as a preprocessing step. The two contexts used for the current work are decided based on the orientations of the torso data and room ID. Room 1 is a classroom with a number of objects (e.g., desks and chairs) in it, and Room 2 is a nearly empty hallway. The dissimilar data distributions can be attributed to the differences in the amount of interruptions in WiFi signals. We followed the same normalization methods as Yang et al. (2023).
@@
ExtraSensory: ExtraSensory is a human activity recognition dataset collected using the ExtraSensory mobile application. A number of features were collected from different cellular devices and smartwatches, though we just used the accelerometer and gyroscope features along with self-reported labels. The contexts are decided based on the location of the phone: hand, pocket, and bag.
@@
Stress Sensing Dataset: This dataset measures the physiological impacts of various kinds of Stress. The dataset is collected using Empatica E4 Wristband to extract features such as EDA (Electrodermal Activity), a skin temperature sensor ($4$ Hz), etc, contributing to a total of 34 features. The data was collected from 30 participants with different demographics, and was labelled ‘Stressed’ or ‘Calm’ based on the stress-inducing or calming tasks. For this work, we collaborated with the authors to identify participants wearing wristbands on both hands for personalization. Activities were annotated as still or moving based on movement patterns, and these annotations will be shared publicly. Contexts are defined by the wristband hand and movement state during data collection.